OpenStudy (anonymous):

3n - 5 = 7n + 11. What is n?

OpenStudy (anonymous):

3n-5=7n+11 -3n -3n _______________ -5 = 4n+11 -11 -11 ______________ -16 = 4n ___ ____ 4 4 -4 = n

OpenStudy (anonymous):

so subtract 3n from both sides then subtract 11 from both sides and you should get -16 = 4n which is n= -4